# FlexTaF: Enhancing Table Reasoning with Flexible Tabular Formats

### v1

@@ 1 Introduction
Table reasoning is a crucial task in natural language processing that aims to automatically extract and infer information from tables (Dong et al. 2022). In this task, a model needs to answer questions based on the information contained in the table, with each question-table pair referred to as an instance. Given the superior commonsense and logical reasoning capabilities of Large Language Models (LLMs), researchers increasingly utilize them for table reasoning, which has become the mainstream method (Chen 2023; Zhang et al. 2024c; Dong and Wang 2024). Therefore, we focus on how to solve the table reasoning task with LLMs in this paper.
@@
Some previous works enhance the table reasoning performance by designing prompts (Cheng et al. 2023; Ye et al. 2023; Zhang et al. 2024e; Lee et al. 2024), such as Chain-of-Table (Wang et al. 2024), which solves table reasoning tasks by prompting the LLM to iteratively reason with natural language and programs. While the above methods provide a fixed tabular format in their prompts, Sui et al. (2024); Singha et al. (2023); Deng et al. (2024) argue that different table reasoning tasks have different most suitable tabular formats. However, existing works primarily analyze from the task aspect without considering that solving different instances demands distinct reasoning skills (Shi et al. 2020; Chen et al. 2023; Cheng et al. 2023; Liu, Wang, and Chen 2024; Chen et al. 2024b). For example, as shown in Figure 1, solving instance (a) involves reasoning with sequential indexing, which is facilitated by the List format. In contrast, solving instance (b) requires identifying columns that satisfy specific conditions, making the Database format more suitable. Therefore, using a fixed tabular format for all instances could limit performance. Additionally, models vary in their reasoning capabilities (Cao et al. 2023; Bhandari et al. 2024; Zhang et al. 2024a), therefore the most suitable tabular format could differ for each model.
@@
To demonstrate the effectiveness of our methods, we conduct experiments on two mainstream table reasoning datasets: WikiTableQuestions (Pasupat and Liang 2015) and TabFact (Chen et al. 2020). Compared to the best performances achieved by using a fixed format with greedy decoding and self-consistency decoding (Wang et al. 2023), FlexTaF-Single and FlexTaF-Vote show average improvements of $2.3\%$ and $4.8\%$ respectively with comparable inference costs, confirming the effectiveness of our methods. Further analysis experiments reveal that some instances can only be correctly solved using a specific format, proving that the most suitable tabular formats for different instances are distinct.
@@ 2 Discussion on Impact of Tabular Formats
We first claim that the most suitable tabular formats for different instances and LLMs are distinct. Since resolving different instances requires diverse abilities (Chen et al. 2023; Cheng et al. 2023; Chen et al. 2024b), it is necessary to tailor tabular formats for each instance accordingly. Additionally, the capabilities of different models vary (Cao et al. 2023; Bhandari et al. 2024; Zhang et al. 2024a), resulting in different tabular formats suitable for different LLMs. We further discuss from the perspective of experimental results in this section.
We select five popular formats, including Markdown, Dict, List, Pandas, and Database formats, following previous works (Singha et al. 2023; Cheng et al. 2023; Wang et al. 2024; Liu, Wang, and Chen 2024). Detailed descriptions of these formats are provided in the supplementary material. Subsequently, we quantitatively analyze how tabular format affects the table reasoning performance of LLMs from the aspects of the instance and the model respectively.
@@ 2.2 The Relationship between Model and Tabular Format
In this subsection, we discuss that different models suit different tabular formats. Table 2 demonstrates that the most suitable tabular format varies across models and the performance gap arises due to different formats. For instance, Llama3 (Meta 2024) exhibits significantly better performance with the Markdown format compared to other formats, while DeepSeek-Coder (Guo et al. 2024) performs better with Dict and Database formats than with Markdown. We employ a Chi-square test to demonstrate significant differences in the distribution of the most suitable tabular formats across different models, as detailed in the supplementary material.
@@ 3.2 FlexTaF-Single
Since there could be multiple formats in $\texttt{F}_{\texttt{M},I}^{*}$, we apply a multi-label classification training method (Godbole and Sarawagi 2004; Tsoumakas and Katakis 2007; Herrera et al. 2016), where each label denotes a tabular format. Moreover, we utilize a binary relevance method (Godbole and Sarawagi 2004) for multi-label classification. Specifically, the tabular formats are serialized and transformed into binary vectors for each instance. During training, we adopt Binary Cross-Entropy Loss, normalized by the number of instances $N$, as follows:
@@ 4.1 Settings
Datasets
We use WikiTableQuestions (Pasupat and Liang 2015) and TabFact (Chen et al. 2020) datasets to evaluate FlexTaF. WikiTableQuestions is a mainstream dataset for the table Question Answering task, containing diverse questions across various domains, which requires answering the question based on the table. TabFact is a prominent dataset for the table fact verification task, which needs to determine whether a claim is entailed or refuted by the table.
@@
Table 2: The accuracy of reasoning using a fixed tabular format with self-consistency decoding (Wang et al. 2023) and FlexTaF-Vote, across four LLMs on WikiTQ and TabFact. The best performance for each LLM and dataset is marked in bold. $\Delta$ denotes the improvement of FlexTaF-Vote relative to the best performance of using a fixed format with self-consistency for each LLM and dataset.
Metric
We employ accuracy as the evaluation metric for WikiTableQuestions and TabFact, following the previous works (Pasupat and Liang 2015; Chen et al. 2020), and use accuracy to evaluate the performance of classification. Accuracy measures the model ability to generate the correct answer, which is achieved only when the predicted result exactly matches the correct answer. Since FlexTaF-Single aims to identify the most suitable format, we adopt accuracy to assess whether the top format predicted by the classifier is in the suitable formats of the instance.
Models
For reasoning, we employ Llama3 (Meta 2024) and DeepSeek-Coder (Guo et al. 2024), and for classification, we use ELECTRA-Large (Clark et al. 2020). Llama3 and DeepSeek-Coder are popular open-source LLMs for their outstanding performance across various tasks. We choose ELECTRA-Large due to its superior performance in language comprehension and question-answering tasks compared to other pre-trained models of similar size (Clark et al. 2020).
Implementation Details
For the WikiTableQuestions and TabFact datasets, we adopt Markdown, Dict, List, Pandas, and Database as tabular formats, which are commonly used in previous works (Singha et al. 2023; Cheng et al. 2023; Ye et al. 2023; Wang et al. 2024) and generally have high performance across datasets with different models. To enhance table reasoning performance, we utilize the Chain-of-Thought (CoT) prompt (Wei et al. 2022) for Markdown and the Program-of-Thought (PoT) prompt (Chen et al. 2023; Gao et al. 2023) for other formats. In addition, we apply 4-shot prompts and 2-shot prompts respectively, since the questions in WikiTableQuestions are more challenging. To improve the performance on WikiTableQuestions, we use different demonstrations for each tabular format, and we explore the performance of FlexTaF with unified demonstrations in §4.3. Detailed prompts are provided in the supplementary material. We train a tabular format classifier for each LLM on each dataset. We provide detailed training information in the supplementary material. For sampling in self-consistency decoding (Wang et al. 2023), we set $temperature=0.1$ which can bring optimal performance across most formats within $temperature\leq 0.8$. To ensure a fair comparison with FlexTaF-Vote, we set $sampling\_n=5$.
@@ 4.2 Main Experiment
Table 1 and Table 2 compare FlexTaF-Single with using a fixed format with greedy decoding, and FlexTaF-Vote with using a fixed format with self-consistency (Wang et al. 2023), respectively. We observe that: (iFlexTaF-Single and FlexTaF-Vote surpass the best results achieved by the fixed format with greedy decoding and self-consistency, by an average of $2.3\%$ and $4.8\%$ respectively, with comparable computational costs. (ii) Compared to flexible tabular formats, the fixed format restricts table reasoning performance even with self-consistency. Moreover, self-consistency with Markdown improves slightly or even decreases compared to greedy decoding. This is attributed to self-consistency instability in CoT prompting (Chen et al. 2024a; Renze and Guven 2024), and diminished instruction following ability at higher sampling temperatures (Zeng et al. 2024; Peeperkorn et al. 2024). Additionally, the tables reveal that:
@@ 4.3 Analysis
We select Llama3-8B for subsequent analytical experiments due to its high reasoning efficiency. Moreover, we conduct most experiments on WikiTQ, because it encompasses more diverse questions (Pasupat and Liang 2015; Chen et al. 2020; Dong et al. 2022; Shi et al. 2020).
Is the impact of tabular formats due to the different demonstrations?
@@
Are the tabular formats suitable for different instances different?
@@
How does the number of candidate tabular formats affect FlexTaF?
@@
We perform experiments using varying numbers of candidate formats and present results in Figure 4, where formats are added in descending order of performance. We observe the following: (i) The performance of FlexTaF-Single gradually stabilizes as the number of candidate formats increases. The classification performance does not always improve due to the increased difficulty with a higher number of labels (Wang et al. 2021; Audibert, Gauffre, and Amini 2024). (iiFlexTaF-Vote varies greatly with the increase in the number of formats due to its reliance on the performance of each format. In particular, FlexTaF-Vote does not exceed FlexTaF-Single when Markdown and Dict are candidates, as it selects from two different answers only by comparing probabilities, which do not accurately indicate correctness (Wang et al. 2023; Portillo Wightman, Delucia, and Dredze 2023; Quevedo et al. 2024). Therefore, we select $5$ formats as candidates in the main experiments, considering the performance of our two methods.
How to further improve the classification performance of FlexTaF-Single?
@@
We analyze both the overall accuracy and the accuracy on instances that can be correctly solved by only one format, as these instances most distinctly highlight the unique features of each format. The classification results are presented in Table 5. We observe that: (i) The classification performance of predicting larger-scale LLMs is better than that of smaller-scale LLMs. This is attributed to the greater robustness of larger-scale LLMs (Howe et al. 2024), which results in more consistent features in instances correctly solved with the same format. (ii) The classification performance of each format is positively correlated with the size of its training data (see the supplementary material). Therefore, future improvements in classification performance can be achieved by reducing noise in the data and increasing the scale of the training data.
How does the data filtering strategy affect FlexTaF-Single?
@@ 4.4 Case Study
To better illustrate that different instances are suitable for different tabular formats, we present an instance from WikiTQ. As illustrated in Figure 6, when utilizing the Dict format, Llama3-70B (Meta 2024) generates an incorrect program that processes all table rows without excluding a special ”-” row. Conversely, when solving the instance with Markdown, the model successfully ignores the ”-” line and correctly identifies the country that won the most gold medals. Therefore, for this instance, the Markdown format proves more suitable than Dict when employing Llama3-70B. Additional instances are provided in the supplementary material.
@@ 5 Related Works
The table reasoning task aims to answer the natural language question based on provided tabular data (Dong et al. 2022; Zhang et al. 2024c; Dong and Wang 2024). LLM-based methods have become predominant in this field due to their superior performance of table understanding, commonsense reasoning, and logical reasoning (Wei et al. 2022; Chen 2023; Zhao et al. 2023; Chen et al. 2023; Gao et al. 2023). Some existing works focus on enhancing table reasoning ability by fine-tuning LLMs (Zhang et al. 2024b; Bian et al. 2024; Zhang et al. 2024d; Patnaik et al. 2024; Wu and Feng 2024; Gardner, Perdomo, and Schmidt 2024; Li et al. 2024). Also, some works improve table reasoning performance by designing prompts (Zhao et al. 2023; Ye et al. 2023; Nahid and Rafiei 2024a, b; Lee et al. 2024; Wang et al. 2024; Zhao et al. 2024) or aggregating diverse results (Ni et al. 2023; Liu, Wang, and Chen 2024). For instance, Mix Self-Consistency (Liu, Wang, and Chen 2024) proposes combining textual and symbolic reasoning results.
@@ B.1 Prompts for Main Experiments
[…]
@@
$\mid 6\mid GaryJones\mid UnitedStates\mid Yamaha\mid 439\mid 0\mid$
$\mid 7\mid JohnDeSoto\mid UnitedStates\mid Suzuki\mid 425\mid 0\mid$
@@
Table 7: The prompt with the Dict format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6. We will release the full prompt upon the acceptance.
@@
    if count == max_riders]
    return countries_with_max_riders[0]
@@
Table 8: The prompt with the List format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6. We will release the full prompt upon the acceptance.
@@
Table 9: The prompt with the Pandas format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6. We will release the full prompt upon the acceptance.
@@
Table 10: The prompt with the Database format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations. We will release the full prompt upon the acceptance.
@@
The prompts for WikiTQ (Pasupat and Liang 2015) in main experiments with a single tabular format are shown in Table 6, Table 7, Table 8, Table 9, and Table 10. And the prompts for TabFact (Chen et al. 2020) in main experiments with a single tabular format are shown in Table 11, Table 12, Table 13, Table 14, and Table 15. It can be found that only the demonstrations with the Database format are different among the prompts for WikiTQ. If we use the same demonstrations as other prompts, the SQL given in the prompt is too complicated to reduce the reasoning performance of the model.

### v2

@@ 1 Introduction
Table reasoning is a crucial task in natural language processing that aims to automatically extract and infer information from tables (Dong et al., 2022). In this task, a model needs to answer questions based on the information contained in the table, with each question-table pair referred to as an instance. Given the superior commonsense and logical reasoning capabilities of Large Language Models (LLMs), researchers increasingly utilize them for table reasoning, which has become the mainstream method (Chen, 2023; Zhang et al., 2024d; Dong & Wang, 2024). Therefore, we focus on how to solve the table reasoning task with LLMs in this paper.
@@
Some previous works enhance the table reasoning performance by designing prompts (Cheng et al., 2023; Ye et al., 2023; Zhang et al., 2024e; Lee et al., 2024), such as Chain-of-Table Wang et al. (2024), which solves table reasoning tasks by prompting the LLM to iteratively reason with natural language and programs. While the above methods provide a fixed tabular format in their prompts, Sui et al. (2024); Singha et al. (2023); Deng et al. (2024) argue that different table reasoning tasks have different most suitable tabular formats. However, existing works primarily analyze from the task aspect without considering that solving different instances demands distinct reasoning skills Shi et al. (2020); Chen et al. (2023); Cheng et al. (2023); Liu et al. (2024a); Chen et al. (2024b). For example, as shown in Figure 1, solving instance (a) involves reasoning with sequential indexing, which is facilitated by the List format. In contrast, solving instance (b) requires identifying columns that satisfy specific conditions, making the Database format more suitable. Therefore, using a fixed tabular format for all instances could limit performance. Additionally, models vary in their reasoning capabilities (Cao et al., 2023; Bhandari et al., 2024; Zhang et al., 2024a), therefore the most suitable tabular format could differ for each model.
@@
To demonstrate the effectiveness of our methods, we conduct experiments on two mainstream table reasoning datasets: WikiTableQuestions (Pasupat & Liang, 2015) and TabFact (Chen et al., 2020). Compared to the best performances achieved by using a fixed format with greedy decoding and self-consistency decoding Wang et al. (2023), FlexTaF-Single and FlexTaF-Vote show average improvements of $2.3\%$ and $4.8\%$ respectively with comparable inference costs, confirming the effectiveness of our methods. Further analysis experiments reveal that some instances can only be correctly solved using a specific format, proving that the most suitable tabular formats for different instances are distinct.
@@ 2 Discussion on Impact of Tabular Formats
We first claim that the most suitable tabular formats for different instances and LLMs are distinct. Since resolving different instances requires diverse abilities Chen et al. (2023); Cheng et al. (2023); Chen et al. (2024b), it is necessary to tailor tabular formats for each instance accordingly. Additionally, the capabilities of different models vary (Cao et al., 2023; Bhandari et al., 2024; Zhang et al., 2024a), resulting in different tabular formats suitable for different LLMs. We further discuss from the perspective of experimental results in this section.
We select five popular formats, including Markdown, Dict, List, Pandas, and Database formats, following previous works Singha et al. (2023); Cheng et al. (2023); Wang et al. (2024); Liu et al. (2024a). Detailed descriptions of these formats are provided in the supplementary material. Subsequently, we quantitatively analyze how tabular format affects the table reasoning performance of LLMs from the aspects of the instance and the model respectively.
@@ 2.2 The Relationship between Model and Tabular Format
In this subsection, we discuss that different models suit different tabular formats. Table 2 demonstrates that the most suitable tabular format varies across models and the performance gap arises due to different formats. For instance, Llama3 Meta (2024) exhibits significantly better performance with the Markdown format compared to other formats, while DeepSeek-Coder Guo et al. (2024) performs better with Dict and Database formats than with Markdown. We employ a Chi-square test to demonstrate significant differences in the distribution of the most suitable tabular formats across different models, as detailed in the supplementary material.
@@ 3.2.1 Classification
Since there could be multiple formats in $\texttt{F}_{\texttt{M},I}^{*}$, we apply a multi-label classification training method (Godbole & Sarawagi, 2004; Tsoumakas & Katakis, 2007; Herrera et al., 2016), where each label denotes a tabular format. Moreover, we utilize a binary relevance method (Godbole & Sarawagi, 2004) for multi-label classification. Specifically, the tabular formats are serialized and transformed into binary vectors for each instance. During training, we adopt Binary Cross-Entropy Loss, normalized by the number of instances $N$, as follows:
@@ 4.1 Settings
4.1.1 Datasets
We use WikiTableQuestions (Pasupat & Liang, 2015) and TabFact (Chen et al., 2020) datasets to evaluate FlexTaF. WikiTableQuestions is a mainstream dataset for the table Question Answering task, containing diverse questions across various domains, which requires answering the question based on the table. TabFact is a prominent dataset for the table fact verification task, which needs to determine whether a claim is entailed or refuted by the table.
@@ 4.1.1 Datasets
Table 2: The accuracy of reasoning using a fixed tabular format with self-consistency decoding Wang et al. (2023) and FlexTaF-Vote, across four LLMs on WikiTQ and TabFact. The best performance for each LLM and dataset is marked in bold. $\Delta$ denotes the improvement of FlexTaF-Vote relative to the best performance of using a fixed format with self-consistency for each LLM and dataset.
4.1.2 Metric
We employ accuracy as the evaluation metric for WikiTableQuestions and TabFact, following the previous works (Pasupat & Liang, 2015; Chen et al., 2020), and use accuracy to evaluate the performance of classification. Accuracy measures the model ability to generate the correct answer, which is achieved only when the predicted result exactly matches the correct answer. Since FlexTaF-Single aims to identify the most suitable format, we adopt accuracy to assess whether the top format predicted by the classifier is in the suitable formats of the instance.
4.1.3 Models
For reasoning, we employ Llama3 (Meta, 2024) and DeepSeek-Coder (Guo et al., 2024), and for classification, we use ELECTRA-Large Clark et al. (2020). Llama3 and DeepSeek-Coder are popular open-source LLMs for their outstanding performance across various tasks. We choose ELECTRA-Large due to its superior performance in language comprehension and question-answering tasks compared to other pre-trained models of similar size (Clark et al., 2020).
4.1.4 Implementation Details
For the WikiTableQuestions and TabFact datasets, we adopt Markdown, Dict, List, Pandas, and Database as tabular formats, which are commonly used in previous works (Singha et al., 2023; Cheng et al., 2023; Ye et al., 2023; Wang et al., 2024) and generally have high performance across datasets with different models. To enhance table reasoning performance, we utilize the Chain-of-Thought (CoT) prompt (Wei et al., 2022) for Markdown and the Program-of-Thought (PoT) prompt (Chen et al., 2023; Gao et al., 2023) for other formats. In addition, we apply 4-shot prompts and 2-shot prompts respectively, since the questions in WikiTableQuestions are more challenging. To improve the performance on WikiTableQuestions, we use different demonstrations for each tabular format, and we explore the performance of FlexTaF with unified demonstrations in §4.3.1. Detailed prompts are provided in the supplementary material. We train a tabular format classifier for each LLM on each dataset. We provide detailed training information in the supplementary material. For sampling in self-consistency decoding Wang et al. (2023), we set $temperature=0.1$ which can bring optimal performance across most formats within $temperature\leq 0.8$. To ensure a fair comparison with FlexTaF-Vote, we set $sampling\_n=5$.
@@ 4.2 Main Experiment
Table 1 and Table 2 compare FlexTaF-Single with using a fixed format with greedy decoding, and FlexTaF-Vote with using a fixed format with self-consistency Wang et al. (2023), respectively. We observe that: (iFlexTaF-Single and FlexTaF-Vote surpass the best results achieved by the fixed format with greedy decoding and self-consistency, by an average of $2.3\%$ and $4.8\%$ respectively, with comparable computational costs. (ii) Compared to flexible tabular formats, the fixed format restricts table reasoning performance even with self-consistency. Moreover, self-consistency with Markdown improves slightly or even decreases compared to greedy decoding. This is attributed to self-consistency instability in CoT prompting Chen et al. (2024a); Renze & Guven (2024), and diminished instruction following ability at higher sampling temperatures Zeng et al. (2024); Peeperkorn et al. (2024). Additionally, the tables reveal that:
@@ 4.3 Analysis
We select Llama3-8B for subsequent analytical experiments due to its high reasoning efficiency. Moreover, we conduct most experiments on WikiTQ, because it encompasses more diverse questions (Pasupat & Liang, 2015; Chen et al., 2020; Dong et al., 2022; Shi et al., 2020).
4.3.1 Is the impact of tabular formats due to the different demonstrations?
@@
4.3.2 Are the tabular formats suitable for different instances different?
@@
4.3.3 How does the number of candidate tabular formats affect FlexTaF?
@@
We perform experiments using varying numbers of candidate formats and present results in Figure 4, where formats are added in descending order of performance. We observe the following: (i) The performance of FlexTaF-Single gradually stabilizes as the number of candidate formats increases. The classification performance does not always improve due to the increased difficulty with a higher number of labels Wang et al. (2021); Audibert et al. (2024). (iiFlexTaF-Vote varies greatly with the increase in the number of formats due to its reliance on the performance of each format. In particular, FlexTaF-Vote does not exceed FlexTaF-Single when Markdown and Dict are candidates, as it selects from two different answers only by comparing probabilities, which do not accurately indicate correctness Wang et al. (2023); Portillo Wightman et al. (2023); Quevedo et al. (2024). Therefore, we select $5$ formats as candidates in the main experiments, considering the performance of our two methods.
4.3.4 How to further improve the classification performance of FlexTaF-Single?
@@
We analyze both the overall accuracy and the accuracy on instances that can be correctly solved by only one format, as these instances most distinctly highlight the unique features of each format. The classification results are presented in Table 5. We observe that: (i) The classification performance of predicting larger-scale LLMs is better than that of smaller-scale LLMs. This is attributed to the greater robustness of larger-scale LLMs Howe et al. (2024), which results in more consistent features in instances correctly solved with the same format. (ii) The classification performance of each format is positively correlated with the size of its training data (see the supplementary material). Therefore, future improvements in classification performance can be achieved by reducing noise in the data and increasing the scale of the training data.
4.3.5 How does the data filtering strategy affect FlexTaF-Single?
@@ 4.4 Case Study
To better illustrate that different instances are suitable for different tabular formats, we present an instance from WikiTQ. As illustrated in Figure 6, when utilizing the Dict format, Llama3-70B Meta (2024) generates an incorrect program that processes all table rows without excluding a special ”-” row. Conversely, when solving the instance with Markdown, the model successfully ignores the ”-” line and correctly identifies the country that won the most gold medals. Therefore, for this instance, the Markdown format proves more suitable than Dict when employing Llama3-70B. Additional instances are provided in the supplementary material.
@@ 5 Related Works
The table reasoning task aims to answer the natural language question based on provided tabular data (Dong et al., 2022; Zhang et al., 2024d; Dong & Wang, 2024). LLM-based methods have become predominant in this field due to their superior performance of table understanding, commonsense reasoning, and logical reasoning (Wei et al., 2022; Chen, 2023; Zhao et al., 2023; Chen et al., 2023; Gao et al., 2023). Some existing works focus on enhancing table reasoning ability by fine-tuning LLMs (Zhang et al., 2024b; Bian et al., 2024; Zhang et al., 2024c; Patnaik et al., 2024; Wu & Feng, 2024; Gardner et al., 2024; Li et al., 2024). Also, some works improve table reasoning performance by designing prompts (Zhao et al., 2023; Ye et al., 2023; Nahid & Rafiei, 2024a; b; Lee et al., 2024; Wang et al., 2024; Zhao et al., 2024) or aggregating diverse results (Ni et al., 2023; Liu et al., 2024a). For instance, Mix Self-Consistency (Liu et al., 2024a) proposes combining textual and symbolic reasoning results.
@@ B.1 Prompts for Main Experiments
$\mid RobinsonR-22\mid Lightutilityhelicopter\mid 1,370lb(635kg)\mid...\mid$
$\mid Bell206B3JetRanger\mid Turboshaftutilityhelicopter\mid 3,200lb(1,451kg)\mid...\mid$
$\mid CH-47DChinook\mid Tandemrotorhelicopter\mid 50,000lb(22,680kg)\mid...\mid$
@@
Table 7: The prompt with the Dict format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
@@
    countries_max = [c for c, count in country_counts.items() if count == max_riders]
    return countries_max[0]
@@
Table 8: The prompt with the List format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
@@
Table 9: The prompt with the Pandas format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
@@
Table 10: The prompt with the Database format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations.
@@
The prompts for WikiTQ Pasupat & Liang (2015) in main experiments with a single tabular format are shown in Table 6, Table 7, Table 8, Table 9, and Table 10. And the prompts for TabFact Chen et al. (2020) in main experiments with a single tabular format are shown in Table 11, Table 12, Table 13, Table 14, and Table 15. It can be found that only the demonstrations with the Database format are different among the prompts for WikiTQ. If we use the same demonstrations as other prompts, the SQL given in the prompt is too complicated to reduce the reasoning performance of the model.